Find the point-slope form of the equation of the line satisfying the given conditions and use this to write the slope-intercept form of the equation.
x-intercept = 3 and y-intercept = 2


Answer by nyc_function(2741)

If the y-intercept is 2, then our point must be (0, 2).

If the x-intercept is 3, then our point must be (3, 0).
We now find the slope.
slope = (0 - 2)/(3 - 0)
slope = -2/3
Plug the slope and one of the points into the point-slope formula.
Here is the point-slope form:
y -(-2) = (-2/3)(x - 0)
To find the slope-intercept form, we solve the point-slope formula above for y.
y -(-2) = (-2/3)(x - 0)
y + 2 = (-2/3)(x) + 0
y = (-2/3)(x) + 0 - 2
y = (-2/3)(x) - 2